OpenStudy (anonymous):

Using elimination solve this. x + y - 4 = 0 2x = 8 + 2y

OpenStudy (anonymous):

I got x=-6 and y = idk

OpenStudy (anonymous):

@HawkCrimson Can u umm help?

OpenStudy (anonymous):

Can you show me how u got it?

OpenStudy (anonymous):

Well idk if i didn't right in the first place. Idk how to find y, which means i more likely got x wrong-_- Sry

OpenStudy (anonymous):

Oh okay then. Thanks for trying =D

OpenStudy (anonymous):

sure im here :)

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

back, ok ill solve it now

OpenStudy (anonymous):

i normally like doing the substitution method but ill do it in elimination for u instead

OpenStudy (anonymous):

Here you can check your answer >> http://calculator.tutorvista.com/math/470/elimination-method-calculator.html

OpenStudy (anonymous):

ik how to do it using the substituting method but im not sure for the eliminating method :/

OpenStudy (anonymous):

Sry tht i mentioned u.

OpenStudy (anonymous):

-_- i take that as a challenge!

OpenStudy (anonymous):

cause i remember that u have to make make the two equations the same then subtract then from each other

OpenStudy (anonymous):

--_--

OpenStudy (anonymous):

Alright just show me the substitution way

OpenStudy (anonymous):

lol ok!

OpenStudy (anonymous):

x = 4 - y 2x = 8 + 2y Therefore, 2(4 - y) = 8 + 2y 8 - 2y = 8 +2y 4y = 0 y = 0 Then, x = 4 - 0 x=4